OpenStudy (ohohaye):

If I have a slope of 5 and points (4, 2) how can I out that into y=mx+b form?

OpenStudy (ohohaye):

I have already graphed it but, I forgot how get it into y=mx+b form

OpenStudy (anonymous):

You can put it into point slope form which is: y – y1 = m(x – x1), then expand it until you reach y=mx+b

OpenStudy (anonymous):

So like: y-2=5(x-4) y-2=5x-20 y=5x-18

OpenStudy (ohohaye):

oh ok, thank you

OpenStudy (anonymous):

You're welcome:)